Ticket Number: HEA-NOC/20041017-2 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20041017 10:28 UTC+1 Problem Start: 20041017 00:00 UTC+1
Ticket Update: 20041018 16:31 UTC+1 Problem End:
The backup link to Geant is down
PRoduction traffic is not affected, but resilience is no longer available
20041017 10:26 BB Investigating
20041017 16:02 BB Visited site - no obvious problems. Will replace cable
between Geant's M40 and HEAnet's Cisco.
20041017 17:47 BB Replaced cable between Geant POP router and HEAnet's
access router, but no change was observed. Have passed diagnostics to Lan
Communications for their analysis.
20041018 10:28 BC Arrange for an onsite ESAT engineer to loop the STM-16
port on phobos. The Rx optical power dropped from -8.8dBm to -32dBm. It
therefore looks as though the Tx laser on the STM-16 card is dead. Asked
Stephen Mulhall of LanComms to arrange for a replacement card to be sent
to City West. Let Geant know that we believe the problem to be with our
20041018 11:32 BC LanComms ticket ref is 2004-SC11249.
20041018 16:16 DC Going onsite at 8 tomorrow morning. Some interruption of
service may be needed when the card is replaced.
This ticket can be monitored at http://www.hea.net/tickets/20041017-2